Filer Manufacturing has 4,294,390 shares of common stock outstanding. The current share price is $59.03, and the book value per share is $5.34. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $55,657,525, has a 0.08 coupon, matures in 13 years and sells for 98 percent of par. The second issue has a face value of $60,631,873, has a 0.08 coupon, matures in 15 years, and sells for 97 percent of par.

The most recent dividend was $2.81 and the dividend growth rate is 0.06. Assume that the overall cost of debt is the weighted average of that implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 0.27.

What is Filer's cost of equity? Enter the answer with 4 decimals (e.g. 0.2345)
Answer #1

Price = 01/(ke-g) 2.81*1.06 D1 = Ke-g= $59.03 6.00% $2.98 0.0505 11.05% 0.1105 Ke = Cost of equity =